# High-Field Microscale NMR Spectroscopy with NV Centers in Dipolarly-Coupled Samples

### v1

@@ I Introduction
During the last decade, the so far most productive leg of quantum technologies Dowling03 , namely, quantum sensing Degen17 , has produced magnetometers able to detect ever weaker fields. Such development has had a deep impact in the realm of nuclear magnetic resonance (NMR) Levitt08 , a field that, despite its unquestionable success, has limitations due to the inherent weakness of target signals necessitates the scanning of millimeter-sized samples. Nitrogen vacancy (NV) centers in diamond Doherty13 , however, reported the detection of signals from smaller samples leading to NMR experiments with unprecedented spatial resolution. NV centers stand out for their capacity to operate at room temperature leading to smaller, and easier to operate magnetometers compared to platforms that require stringent conditions such as, e.g., superconducting quantum interference devices (SQUIDs).
@@
NMR spectroscopy reveals frequency shifts relative to a base Larmor frequency which encode structural information about sample molecules as well as of their surrounding environment. In this scenario, one of the most impressive results produced by NV based NMR sensors –enabled by heterodyne protocols that overcome the resolution boundary posed by the coherence time of the NV centers Boss17 ; Schmitt17 – is the record of spectral features from picoliter volume samples with high-resolution Glenn18 . Nevertheless, in thermally polarized samples, these protocols require a high nuclear spin concentration (typically, highly protonated samples) and a significant number of repetitions to obtain meaningful results. A possibility for detecting samples at lower concentrations and achieving more competitive protocols is to hyperpolarize the samples in a previous step Arunkumar21 ; Bucher20 . In addition, performing the experiment at high fields directly provides higher polarization rates while it facilitates the extraction of relevant information from the recorded spectra, as chemical shifts increase and J-couplings become clearer Ugurbil03 . Although NV-NMR spectroscopy at high fields has not yet been experimentally demonstrated, recent proposals have introduced protocols that enable the acquisition of high-resolution spectra in strong external magnetic fields Munuera23 ; Meinel23 ; Alsina23 . In particular, our proposal AERIS Munuera23 operates encoding the target nuclear energy shifts in the amplitude variation of the sample’s longitudinal magnetization which oscillates at a tunable rate –i.e., at a slow rate of, typically, tens of KHz even at large fields– during consecutive detections. Recent publications indicate that the experimental implementation of AERIS is just a matter of time Daly23 .
An important limitation of state-of-the-art NV-NMR spectroscopy is caused by the strong dipolar coupling among target nuclei, which results in intricate spectra that challenge data interpretation. This becomes especially critical in solid-state material research, where homonuclear dipole-dipole interactions hinder subtler couplings (heteronuclear interactions are less challenging as they can be eliminated by driving only one of the species). The NMR community has dedicated significant efforts to mitigate the impact of dipolar interactions Madhu16 , and today solid-state NMR spectroscopy is extensively used in distinct research areas: In pharmaceutics it characterizes active pharmaceutical ingredients (APIs) and their interaction with excipients (inactive substances added to a drug that serve various purposes such as binding or preserving the API) Pisklak14 ; Pisklak16 ; Nie16 , among many other applications (see Marchetti21 for an extensive review on the topic). In epidemiology, it provides key insights of the structure of molecules related with diseases as present in our societies as Alzeimer Tycko11 . In energy storage research is used to characterize the local structure of solid materials used in batteries and fuel cells Pecher17 .
These areas, along with many others, would largely benefit form sensors able to produce narrower spectral lines from smaller solid-state samples. NV-NMR spectroscopy emerges as the prominent technique to access samples in the microscale regime, however, it remains limited to the liquid state scenario where the dipole-dipole interactions get naturally averaged out due to fast molecular motion, leaving a plethora of relevant applications out of its range of action.
@@
Figure 1: (a) Relative positions of axis $A$ (clear-blue), $\bar{A}$ (dark-blue), $B$ (clear-magenta), $\bar{B}$ (dark-magenta). The motion of the magnetization inbetween LG4 blocks ($A\bar{A}\bar{B}B$) is shown in purple. This evolution is described as a rotation in the plane (in yellow) perpendicular to the $C$ axis in accordance with the effective Hamiltonian in Eq. (4) for a single $\delta_{i}^{*}$. (b) (Top) Magnetization rotations during each of the four RF drivings. (Bottom) Projection of the magnetization onto $\hat{z}$ as it rotates during a full LG4 block. This projection determines the magnetic signal for the NV ensemble sensor.
@@
In this work, we devise a protocol that overcomes these limitations and enables high resolution NV-NMR spectroscopy of ordered samples with strong homonuclear dipolar coupling at elevated external magnetic fields, allowing to take advantage from the higher polarization rates and stronger chemical shifts. Our protocol features the delivery of two radiation channels –radio-frequency (RF) and microwave (MW)– synchronized with measurements on an NV ensemble magnetometer. The RF channel drives the sample with a twofold purpose. On the one hand it effectively decouples the target nuclear spins, diminishing the effect of strong homonuclear couplings in the recorded spectra, and enabling the obtention of nuclear energy shifts. Remarkably, the decoupling benefits from increasing RF intensities could be specially effective in the small volume regime, where current RF antennas have reported nuclear spin rotation velocities within the range of tens to hundreds of kHz Herb20 ; Yudilevich23 . On the other hand, the RF bridges the interaction among NV sensors and fast rotating nuclear spin by generating a slow-frequency NMR signal trackable by the NV sensor. Simultaneously, the MW channel delivers a tailored pulse sequence to the NV ensemble enabling the detection of the magnetic field emitted by the driven sample. This sequence is interspersed with measurements of the sensor’s state to construct the spectra in a heterodyne frame leading to a spectrum only limited by the nuclear sample coherence. Finally, we provide analytical expressions that map the detected resonances with target energy shifts.
II The Method
II.1 RF modulation of nuclear spins
@@
Lee and Goldburg (LG) showed in a seminal paper Lee65 that an off-resonant continuous RF field cancels, up to first order, the contribution to the nuclear spin dynamics of homonuclear dipole-dipole interactions if the LG condition $\Delta=\pm\Omega/\sqrt{2}$ holds. Here, $\Delta=\omega_{L}-\omega_{d}$ is the detuning between the carrier frequency of the RF driving field ($\omega_{d}$) and the Larmor precession of the spins ($\omega_{L}$), and $\Omega$ is the Rabi frequency of the RF driving. Subjecting a spin ensemble to an off-resonant RF field leads to collective nuclear spin rotations along an axis tilted with respect to $\hat{z}$ (the direction of the static magnetic field). More specifically, the tilted axis –in the following $P$– has a component $\frac{\Delta}{\sqrt{\Omega^{2}+\Delta^{2}}}$ along $\hat{z}$, while its projection on the orthogonal $xy$ plane is $\frac{\Omega}{\sqrt{\Omega^{2}+\Delta^{2}}}$.
@@
Further developments have built upon the original LG sequence demonstrating the ability to remove higher order contributions of the dipole-dipole interaction, thus leading to even narrower spectral lines. Prominent examples are the frequency-switched (FSLG) Waugh72 and phase-modulated (PMLG) Vinogradov99 versions of the original LG sequence. Our protocol incorporates the advanced LG4 sequence Halse13 over nuclei, which exhibits remarkable decoupling rates and enhanced robustness against RF control errors. The LG4 consists on concatenated blocks of four consecutive off-resonant RF drivings, all complying with the LG condition, leading to rotations along four different axes. This is, the rotation axis $P$ alternates among $A,\bar{A},\bar{B}$ and $B$, whose relative positions are illustrated in Fig. 1 (a). Note that, at each block, nuclear spins undergo two sets of complementary rotations along axis pointing in opposite directions ($A,\bar{A}$ and $B,\bar{B}$).
@@
The nuclear spin Hamiltonian during each individual rotation of the LG4 sequence reads (see Appendix A)
@@
where $\delta_{i}$ is the target nuclear shift of the $i$th spin (its sign depends on the direction of the rotation, positive value ``$+\delta_{i}"$ is assigned to rotations along $A$ and $B$, and the negative value ``$-\delta_{i}"$ to rotation along $\bar{A}$ and $\bar{B}$), $\bar{\Omega}=\sqrt{\Delta^{2}+\Omega^{2}}$ is the effective Rabi frequency, and the spin operator $I^{i}_{P}$ takes one of the following forms
@@
According to the LG4 scheme Halse13 , the phase of the driving is set to $\alpha=55^{\circ}$ to minimize the line-width of the resonances.
Note that in Eq. (1) we assume that the internuclear interaction Hamiltonian $H_{\rm nn}=\sum_{i>j}^{N}\frac{\mu_{0}\gamma^{2}_{n}\hbar}{4\pi r_{i,j}^{3}}\bigg{[}\vec{I}_{i}\cdot\vec{I}_{j}-3(\vec{I}_{j}\cdot\hat{r}_{i,k})(\vec{I}_{j}\cdot\hat{r}_{i,j})\bigg{]}$ can be neglected due to the introduced decoupling sequence. This assumption simplifies the subsequent analysis. However, $H_{\rm nn}$ will be taken into account in the numerical model in Sec. III.
@@
Hereafter, we often refer to $s(t)$ as the signal, as it constitutes the target field for the NV ensemble sensor. In fact, its amplitude $\Gamma$, phase $\phi$, and static bias $b$ depend on the configuration of the nuclear spin ensemble and thereby on the $\delta_{i}$ energy shifts (see Appendix C), so detecting and properly reading $s(t)$ enables to unravel the desired information.
Consequently, the LG4 meets a twofold goal. Namely: (i) It results in a nuclear spin dynamics with minimal effect from the dipole-dipole interaction (see Appendix A for the full derivation of the Hamiltonian in Eq. (1)), enabling the identification of the weaker but interesting $\delta_{i}$ shifts. (ii) It induces a tunable rotation speed in the sample ($\propto\bar{\Omega}$, see Eq. (3)), facilitating the interaction between nuclear spins and the NV ensemble sensor even at high external magnetic fields. Regarding point (ii), it is important to note that without using RF drivings on the sample, standard techniques based on imprinting in the NVs a rotation speed comparable to the nuclear Larmor frequency would necessitate the application of unrealistic MW fields. For context, in a magnetic field of approximately 2.35 Tesla, hydrogen spins rotate at a speed of $(2\pi)\times 100$ MHz, producing a signal hardly trackable by an NV ensemble sensor operating with conventional methods. Glenn18 ; Bucher20 ; Arunkumar21 .
@@
Interestingly, with this RF control, the nuclear spins governed by Eq. (1) would perform a complete turn at each RF driving, constantly returning to their initial configuration if it were not for the $\delta_{i}$ shifts. These shifts slightly alter the nuclear spin state (i.e., the sample magnetization), thus imprinting a slower motion within the sample. More specifically, the sample magnetization at the end of each LG4 block is determined by a set of energy shifts $\delta^{*}_{j}$ (distinct from $\delta_{i}$) according to the effective Hamiltonian (for details see Appendix A):
@@
II.2 Harvesting nuclear spin parameters with the NV ensemble
II.2.1 Geometrical interpretation of the phase accumulation
The target magnetic field over the NV ensemble sensor is a concatenation of the sinusoidal signals in Eq. (3) (see lower panel in Fig. 1(b)). A particular RF field at the $k^{\text{th}}$ LG4 block (note that, the accumulative character of the rotations imposed by Eq. (4) make it crucial to identify the number of the block from now on), produces a nuclear spin rotation around a certain axis ($A$, $\bar{A}$, $B$ or $\bar{B}$) where the amplitude $\Gamma_{k}$ of the resulting signal $s_{k}(t)=\Gamma_{k}\cos{(\bar{\Omega}t+\phi_{k})}+b_{k}$ is directly proportional to $\vec{M}_{k}^{\perp}$ (i.e., to the magnetization component which is orthogonal to the rotation axis –$A$, $\bar{A}$, $B$ or $\bar{B}$– at the start of each RF driving), and the phase $\phi_{k}$ corresponds to the angle between $\vec{M}_{k}^{\perp}$ and $\hat{z}^{\perp}$. The latter is the component of $\hat{z}$ that lies on the plane perpendicular to the rotation axis. See the lower panel in Fig. 2 (a) and Appendix C for more details.
We now use this geometric description to analyze the phase accumulated by each NV in the ensemble sensor when subjected to a generic pulse sequence. For this analysis, we choose a standard Carr-Purcell-Meiboom-Gill (CPMG) sequence Purcell46 ; Meiboom58 . In order to keep the discussion accessible, we focus on the signal produced by the nuclear spins rotating around $A$ and limit ourselves to an scenario involving a single effective energy shift, $\delta_{i}^{*}$. Note, however, that the following results and the consequent conclusions are valid for signals produced by nuclear spins rotating around axes $B$, $\bar{A}$ and $\bar{B}$ and in situations involving multiple shifts.
The phase accumulated by an NV center interacting with the signal $s_{k}(t)$ and subjected to the CPMG sequence reads (see Appendix B)
@@ II.2.1 Geometrical interpretation of the phase accumulation
Figure 2: (a) Illustration of a CPMG pulse block (upper panel) and its geometric interpretation (lower panel). This panel shows a projection of the sphere in Fig. 1 (b) viewed in a direction parallel to axis $A$. This view facilitates the representation of the projections onto the plane perpendicular to A of (i) The magnetization vector, denoted as $\vec{M}_{k}^{\perp}$, and (ii) The $\hat{z}$ axis, referred to as $\hat{z}^{\perp}$. In addition, it shows the trajectory followed by a magnetization vector during a rotation around $A$ (purple circle) and after successive LG4 blocks (yellow ellipse). (b) Upper panel, pulse block of our tailored sequence where the initial $\pi$ pulse is delivered at a time $t_{1}$. With this control, the phase accumulation of the NV is proportional to the projection of $\vec{M}_{k}^{\perp}$ onto $\hat{l}$ (shown in red), an axis tilted away from $\hat{z}^{\perp}$ and aligned with the major axis of the yellow ellipse for optimal contrast. (c) Evolution of the projection of $\vec{M}_{k}^{\perp}$ onto axis $\hat{z}^{\perp}$ (red) and onto axis $\hat{l}$ (blue). The amplitude of the projection onto $\hat{l}$, resulting from the sequence in panel (b), reaches the maximum value of 1. As the phase accumulated by the NV is directly proportional to this projection, the timing of the pulses in (b) ensures the maximum phase accumulation amplitude.
@@
where $\rho_{i}$ is the spin density of the $i$th nucleus, $D_{0}$ is detemined by the pulse sequence and $\nu_{0}$ depends on both the pulse sequence and the initial nuclear state. A formal derivation of (7) as well as further details can be found in Appendix C. The 00 subindex in Eq. (7) indicate that all parameters correspond to the reference CPMG sequence (note that, in the next section we derive an improved sequence). Thus, the NV response enclosed in Eq. (7) consists on a sum of sinusoidal functions that encode the different $\delta_{i}^{*}$ which can be then extracted via standard Fourier transform. Finally, $\delta_{i}$ targets can be obtained via a direct application of Eq. (5).
II.2.2 Sensing MW pulse sequence
@@ III Results
We test our protocol by simulating its implementation to detect the chemical shifts of hydrogen nuclear spins in ethanol molecules (${\rm C_{2}H_{6}O}$) at high external magnetic field. Although ethanol typically exists as a liquid, we employ its solid configuration as an example of an ordered sample with strong homonuclear dipole-dipole couplings. In particular, ethanol molecules exhibit dipolar couplings of up to 17 kHz (the proton attached to the oxygen shows limited dipole interaction with the rest of the system so we exclude it from the simulations). High external fields improve nuclear magnetic resonance procedures, not only because it yields higher polarization rates, but also because it enhances the weaker, and thus harder to detect, energy shifts. Here we consider $B_{0}=2.1$ T and chemical shifts of 3.66 ppm and 1.19 ppm, resulting in shifts of approximately 327 Hz and 106 Hz, respectively.

### v2

# Pulse Sequence Design for High Field NMR with NV Centers in Dipolarly Coupled Samples

@@ Introduction
Over the past decade, quantum sensing–a notably prolific branch of quantum technologies [1, 2], has produced magnetometers able to detect ever weaker fields. Such development has had a deep impact in the realm of nuclear magnetic resonance (NMR) [3], a field that, despite its unquestionable success, has limitations due to the inherent weakness of target signals necessitates the scanning of millimeter-sized samples. Nitrogen vacancy (NV) centers in diamond [4], however, reported the detection of signals from smaller samples leading to NMR experiments with unprecedented spatial resolution. NV centers stand out for their capacity to operate at room temperature leading to smaller, and easier to operate magnetometers compared to platforms that require stringent conditions such as, e.g., superconducting quantum interference devices (SQUIDs).
@@
NMR spectroscopy reveals frequency shifts relative to a base Larmor frequency which encode structural information about sample molecules as well as of their surrounding environment. In this scenario, one of the most impressive results produced by NV based NMR sensors –enabled by heterodyne protocols that overcome the resolution boundary posed by the coherence time of the NV centers [5, 6]– is the record of spectral features from picoliter volume samples with high-resolution [1]. Nevertheless, in thermally polarized samples, these protocols require a high nuclear spin concentration (typically, highly protonated samples) and a significant number of repetitions to obtain meaningful results. A possibility for detecting samples at lower concentrations and achieving more competitive protocols is to hyperpolarize the samples in a previous step [8, 9]. In addition, performing the experiment at high fields directly provides higher polarization rates while it facilitates the extraction of relevant information from the recorded spectra, as chemical shifts increase and J-couplings become clearer [10]. Although NV-NMR spectroscopy at high fields has not yet been experimentally demonstrated, recent proposals have introduced protocols that enable the acquisition of high-resolution spectra in strong external magnetic fields [2, 13, 11]. In particular, our proposal AERIS [2] operates encoding the target nuclear energy shifts in the amplitude variation of the sample’s longitudinal magnetization which oscillates at a tunable rate –i.e., at a slow rate of, typically, tens of KHz even at large fields– during consecutive detections.
@@
Figure 1: (a) Relative positions of axis AA (clear-blue), $\bar{A}$ (dark-blue), $B$ (clear-magenta), $\bar{B}$ (dark-magenta). The motion of the magnetization inbetween LG4 blocks ($A\bar{A}\bar{B}B$) is shown in purple. This evolution is described as a rotation in the plane (in yellow) perpendicular to the $C$ axis in accordance with the effective Hamiltonian in Eq. (4) for a single $\delta_{i}^{*}$. (b) (Top) Magnetization rotations during each of the four RF drivings. (Bottom) Projection of the magnetization onto $\hat{z}$ as it rotates during a full LG4 block. This projection determines the magnetic signal for the NV ensemble sensor.
An important limitation of state-of-the-art NV-NMR spectroscopy is caused by the strong dipolar coupling among target nuclei, which results in intricate spectra that challenge data interpretation. This becomes especially critical in solid-state material research, where homonuclear dipole-dipole interactions hinder subtler couplings (heteronuclear interactions are less challenging as they can be eliminated by driving only one of the species). The NMR community has dedicated significant efforts to mitigate the impact of dipolar interactions [15], and today solid-state NMR spectroscopy is extensively used in distinct research areas: In pharmaceutics it characterizes active pharmaceutical ingredients (APIs) and their interaction with excipients (inactive substances added to a drug that serve various purposes such as binding or preserving the API) [16, 17, 18], among many other applications (see [19] for an extensive review on the topic). In epidemiology, it provides key insights of the structure of molecules related with diseases as present in our societies as Alzheimer [20]. In energy storage research is used to characterize the local structure of solid materials used in batteries and fuel cells [21].
These areas, along with many others, would largely benefit from sensors able to produce narrower spectral lines from smaller solid-state samples, especially over material surfaces where conventional NMR techniques are highly constrained. While NV-NMR spectroscopy emerges as the prominent technique to access samples in the microscale regime, it is currently limited to liquid state samples where dipole-dipole interactions get naturally averaged out due to fast molecular motion, leaving a plethora of solid-state applications out of its range of action. Extending high-field NV-NMR protocols, such as AERIS, to solid-state samples by effectively decoupling dipolar interactions, would unlock these applications [22, 23].
@@
In this work, we devise a protocol that overcomes these limitations and enables high resolution NV-NMR spectroscopy of single crystals with strong homonuclear dipolar coupling at elevated external magnetic fields, allowing to take advantage from the higher polarization rates and stronger chemical shifts. Our protocol features the delivery of two radiation channels –radio-frequency (RF) and microwave (MW)– synchronized with measurements on an NV ensemble magnetometer. The RF channel drives the sample with a twofold purpose. On the one hand it effectively decouples the target nuclear spins, diminishing the effect of strong homonuclear couplings in the recorded spectra, and enabling the obtention of nuclear energy shifts. Remarkably, the decoupling benefits from increasing RF intensities could be specially effective in the small volume regime, where the driving field tends to be more homogeneous. In this regime, current RF antenna designs have demonstrated nuclear spin rotation rates in the tens to hundreds of kilohertz range [24, 25]. On the other hand, the RF bridges the interaction among NV sensors and fast rotating nuclear spin by generating a slow-frequency NMR signal trackable by the NV sensor. Simultaneously, the MW channel delivers a tailored pulse sequence to the NV ensemble enabling the detection of the magnetic field emitted by the driven sample. This sequence is interspersed with measurements of the sensor’s state to construct the spectra in a heterodyne frame leading to a spectrum only limited by the nuclear sample coherence. Finally, we provide analytical expressions that map the detected resonances with target energy shifts.
Methods
RF modulation of nuclear spins
@@ Methods
Lee and Goldburg (LG) showed in a seminal paper [26] that an off-resonant continuous RF field cancels, up to first order, the contribution to the nuclear spin dynamics of homonuclear dipole-dipole interactions if the LG condition $\Delta=\pm\Omega/\sqrt{2}$ holds. Here, $\Delta=\omega_{L}-\omega_{d}$ is the detuning between the carrier frequency of the RF driving field ($\omega_{d}$) and the Larmor precession of the spins ($\omega_{L}$), and $\Omega$ is the Rabi frequency of the RF driving. Subjecting a spin ensemble to an off-resonant RF field leads to collective nuclear spin rotations along an axis tilted with respect to $\hat{z}$ (the direction of the static magnetic field). More specifically, the tilted axis –in the following PP– has a component $\frac{\Delta}{\sqrt{\Omega^{2}+\Delta^{2}}}$ along $\hat{z}$, while its projection on the orthogonal $xy$ plane is $\frac{\Omega}{\sqrt{\Omega^{2}+\Delta^{2}}}$.
@@
Further developments have built upon the original LG sequence demonstrating the ability to remove higher order contributions of the dipole-dipole interaction, thus leading to even narrower spectral lines. Prominent examples are the frequency-switched (FSLG) [28] and phase-modulated (PMLG) [29] versions of the original LG sequence. Our protocol incorporates the advanced LG4 sequence [30] over nuclei, which exhibits remarkable decoupling rates and enhanced robustness against RF control errors. The LG4 consists on concatenated blocks of four consecutive off-resonant RF drivings, all complying with the LG condition, leading to rotations along four different axes. This is, the rotation axis $P$ alternates among $A,\bar{A},\bar{B}$ and $B$, whose relative positions are illustrated in Fig. 1 (a). Note that, at each block, nuclear spins undergo two sets of complementary rotations along axis pointing in opposite directions ($A,\bar{A}$ and $B,\bar{B}$). To further illustrate the journey of the magnetization during an LG4 block we include an animation [31].
@@
The nuclear spin Hamiltonian during each individual rotation of the LG4 sequence reads (see Supplementary Information [27])
@@
where $\delta_{i}$ is the target nuclear shift of the $i$th spin (its sign depends on the direction of the rotation, positive value $\delta_{i}$" is assigned to rotations along $A$ and $B$, and the negative value $\delta_{i}$" to rotation along $\bar{A}$ and $\bar{B}$), $\bar{\Omega}=\sqrt{\Delta^{2}+\Omega^{2}}$ is the effective Rabi frequency, and the spin operator $I^{i}_{P}$ takes one of the following forms
@@
According to the LG4 scheme [30], the phase of the driving is set to $\alpha=55^{\circ}$ to minimize the line-width of the resonances.
Note that in Eq. (1) we assume that the internuclear interaction Hamiltonian $H_{\rm nn}=\sum_{i>j}^{N}\frac{\mu_{0}\gamma^{2}_{n}\hbar}{4\pi r_{i,j}^{3}}\bigg{[}\vec{I}_{i}\cdot\vec{I}_{j}-3(\vec{I}_{j}\cdot\hat{r}_{i,k})(\vec{I}_{j}\cdot\hat{r}_{i,j})\bigg{]}$ can be neglected due to the introduced decoupling sequence. This assumption simplifies the subsequent analysis. However, $H_{\rm nn}$ will be taken into account in the numerical model in the results section.
@@
Hereafter, we often refer to $s(t)$ as the signal, as it constitutes the target field for the NV ensemble sensor. In fact, its amplitude $\Gamma$, phase $\phi$, and static bias $b$ depend on the configuration of the nuclear spin ensemble and thereby on the $\delta_{i}$ energy shifts (see [27]), so detecting and properly reading $s(t)$ enables to unravel the desired information.
Consequently, the LG4 meets a twofold goal. Namely: (i) It results in a nuclear spin dynamics with minimal effect from the dipole-dipole interaction (see [27]for the full derivation of the Hamiltonian in Eq. (1)), enabling the identification of the weaker but interesting $\delta_{i}$ shifts. (ii) It induces a tunable rotation speed in the sample ($\propto\bar{\Omega}$, see Eq. (3)), facilitating the interaction between nuclear spins and the NV ensemble sensor even at high external magnetic fields. Regarding point (ii), it is important to note that without using RF drivings on the sample, standard techniques based on imprinting in the NVs a rotation speed comparable to the nuclear Larmor frequency would necessitate the application of unrealistic MW fields. For context, in a magnetic field of approximately 2.35 Tesla, hydrogen spins rotate at a speed of $(2\pi)\times 100$ MHz, producing a signal hardly trackable by an NV ensemble sensor operating with conventional methods [1, 9, 8].
@@
Interestingly, with this RF control, the nuclear spins governed by Eq. (1) would perform a complete turn at each RF driving, constantly returning to their initial configuration if it were not for the $\delta_{i}$ shifts. These shifts slightly alter the nuclear spin state (i.e., the sample magnetization), thus imprinting a slower motion within the sample. More specifically, the sample magnetization at the end of each LG4 block is determined by a set of energy shifts $\delta^{*}_{j}$ (distinct from $\delta_{i}$) according to the effective Hamiltonian:
@@
Harvesting nuclear spin parameters with the NV ensemble
Geometrical interpretation of the phase accumulation
Figure 2: (a) Illustration of a CPMG pulse block (upper panel) and its geometric interpretation (lower panel). This panel shows a projection of the sphere in Fig. 1 (b) viewed in a direction parallel to axis $A$, as represented with the eye symbol. This view facilitates the representation of the projections onto the plane perpendicular to A of (i) The magnetization vector, denoted as $\vec{M}_{k}^{\perp}$, and (ii) The $\hat{z}$ axis, referred to as $\hat{z}^{\perp}$. In addition, it shows the trajectory followed by a magnetization vector during a rotation around AA (purple circle) and after successive LG4 blocks (yellow ellipse). (b) Upper panel, pulse block of our tailored sequence where the initial $\pi$ pulse is delivered at a time $t_{1}$. With this control, the phase accumulation of the NV is proportional to the projection of $\vec{M}_{k}^{\perp}$ onto $\hat{l}$ (shown in red), an axis tilted away from $\hat{z}^{\perp}$ and aligned with the major axis of the yellow ellipse for optimal contrast. (c) Evolution of the projection of $\vec{M}_{k}^{\perp}$ onto axis $\hat{z}^{\perp}$ (red) and onto axis $\hat{l}$ (blue). The amplitude of the projection onto $\hat{l}$, resulting from the sequence in panel (b), reaches the maximum value of 1. As the phase accumulated by the NV is directly proportional to this projection, the timing of the pulses in (b) ensures the maximum phase accumulation amplitude.
The target magnetic field over the NV ensemble sensor is a concatenation of the sinusoidal signals in Eq. (3) (see lower panel in Fig. 1(b)). A particular RF field at the $k^{\text{th}}$ LG4 block (note that, the accumulative character of the rotations imposed by Eq. (4) make it crucial to identify the number of the block from now on), produces a nuclear spin rotation around a certain axis ($A$, $\bar{A}$, $B$ or $\bar{B}$) where the amplitude $\Gamma_{k}$ of the resulting signal $s_{k}(t)=\Gamma_{k}\cos{(\bar{\Omega}t+\phi_{k})}+b_{k}$ is directly proportional to $\vec{M}_{k}^{\perp}$ (i.e., to the magnetization component which is orthogonal to the rotation axis –$A$, $\bar{A}$, $B$ or $\bar{B}$– at the start of each RF driving), and the phase $\phi_{k}$ corresponds to the angle between $\vec{M}_{k}^{\perp}$ and $\hat{z}^{\perp}$. The latter is the component of $\hat{z}$ that lies on the plane perpendicular to the rotation axis. See the lower panel in Fig. 2 (a) and [27] for more details.
We now use this geometric description to analyze the phase accumulated by each NV in the ensemble sensor when subjected to a generic pulse sequence. For this analysis, we choose a standard Carr-Purcell-Meiboom-Gill (CPMG) sequence [33, 34]. In order to keep the discussion accessible, we focus on the signal produced by the nuclear spins rotating around $A$ and limit ourselves to an scenario involving a single effective energy shift, $\delta_{i}^{*}$. Note, however, that the following results and the consequent conclusions are valid for signals produced by nuclear spins rotating around axes $B$, $\bar{A}$ and $\bar{B}$ and in situations involving multiple shifts.
The phase accumulated by an NV center interacting with the signal $s_{k}(t)$ and subjected to the CPMG sequence reads (see [27])
@@
where $\rho_{i}$ is the spin density of the $i$th nucleus, $D_{0}$ is detemined by the pulse sequence and $\nu_{0}$ depends on both the pulse sequence and the initial nuclear state. A formal derivation of (7) as well as further details can be found in [27]. The 0 subindex in Eq. (7) indicate that all parameters correspond to the reference CPMG sequence (note that, in the next section we derive an improved sequence). Thus, the NV response enclosed in Eq. (7) consists on a sum of sinusoidal functions that encode the different $\delta_{i}^{*}$ which can be then extracted via standard Fourier transform. Finally, $\delta_{i}$ targets can be obtained via a direct application of Eq. (5).
Sensing MW pulse sequence
@@
Results
We test our protocol by simulating its implementation to detect the chemical shifts of hydrogen nuclear spins in ethanol molecules ($\rm C_{2}H_{6}O$}) at high external magnetic field. Although ethanol typically exists as a liquid, we employ its solid configuration as an example of an ordered sample with strong homonuclear dipole-dipole couplings. In particular, ethanol molecules exhibit dipolar couplings of up to 17 kHz (the proton attached to the oxygen shows limited dipole interaction with the rest of the system so we exclude it from the simulations). High external fields improve nuclear magnetic resonance procedures, not only because it yields higher polarization rates, but also because it enhances the weaker, and thus harder to detect, energy shifts. In our simulations, we consider an external field of $B_{0}=2.1$ T, with chemical shifts of 3.66 ppm for three of the protons and 1.19 ppm for the other two, corresponding to frequency shifts of approximately 327 Hz and 106 Hz, respectively.